OpenStudy (anonymous):

What is the length of the hypotenuse of the right triangle shown below? If necessary, round your answer to the nearest hundredth.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

can you determine the lengths of the sides?

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

the lengths of the sides means how long each side is

OpenStudy (anonymous):

i'm confused on how to get the answer

OpenStudy (anonymous):

do you know how to determine a length when you have two points ?

OpenStudy (anonymous):

not one bit,

OpenStudy (anonymous):

there are a few ways to do this. but the easiest is just use the distance equation \[distance = \sqrt{(x_{2} - x_{1})^{2} + (y_{2} - y_{1})^{2}}\] the points on the hypotenuse of that triangle are (4, 2) and (-5, -3) it looks like so just calculate the distance between those two points

OpenStudy (anonymous):

are you in conexus

OpenStudy (anonymous):

that's the easy way. the other way is to calculate the lengths of the sides separately just by looking at the picture, the length of the top side is 9 and the length of the left side is 5 9^2 + 5^2 = c^2

OpenStudy (anonymous):

i don't know what conexus is but it called graphing data for my work?

OpenStudy (anonymous):

 maybe that will help ?